# Dark Matter (H)eats Young Planets

### v2

@@ Abstract
For comparison we show current direct detection constraints as gray contours. Here we show the maximal reach in the sensitivity to the direct nuclear recoil process, which are least model dependent. The reach is dominated by the CRESST results at lower masses for spin-independent Angloher et al. (2023), and spin-dependent Angloher et al. (2022) cross sections, while at higher masses the LUX Akerib et al. (2017) results dominate for spin-dependent interactions, while the XENON experiment dominates the spin-independent sensitivity Aprile et al. (2023). For large spin-independent cross sections, we also show the XQC constraint Erickcek et al. (2007); Mahdawi and Farrar (2018). Note that searches based on the Migdal effect Zhang et al. (2022) could potentially reach lower DM masses, however, the results have large theoretical uncertainties Cox et al. (2023), and are not displayed here.
@@
The sensitivity ceiling of direct detection experiments at large cross sections for light DM from Ref. Cappiello (2023) and for heavy DM from Refs. Kavanagh (2018); Cappiello et al. (2021) is shown. Since at the large cross sections considered the scaling with atomic mass number is uncertain Digman et al. (2019), we do not rescale the results for spin-dependent interactions and a more detailed analysis to obtain a precise ceiling values is needed. Note that in contrast the sensitivity of our constraints and suggested search is only affected at cross sections above $10^{-22}\text{cm}^{2}$, where large drift times would require a more detailed analysis of the DM distribution Acevedo et al. (2021); Leane and Smirnov (2022) and annihilation outside the core.
@@
We also show complementary cosmology-dependent constraints from the CMB Boddy and Gluscevic (2018). DM which scatters with a velocity-independent cross-section can also be constrained from the lack of MW satellites produced in such models Nadler et al. (2019). However, as the DM scattering rates could be different in the early universe Elor et al. (2023), those bounds are not as universal as the shown late-time sensitivity.
@@ Discussion
In this paper, we have studied the inhibition of the formation of Jupiter in our solar system and Jovian planets in the Milky Way galaxy by the capture and annihilation of DM. We have shown that the heat generated by DM annihilation can increase the temperature of the protoplanet above the temperature at which the gas mixture evaporates, and determined the corresponding constraints on the parameter space of spin-(in)dependent DM. These constraints are competitive for relatively strongly interacting DM candidates, above the “ceiling” of direct detection constraints. We have shown that these constraints depend on the ambient temperature at formation, with stronger effects for planets forming further away from the star.
We have also estimated the projected sensitivities from the existence of Jovian planets in the GC, where the DM density is higher and the typical velocity dispersion lower. Future expolanet observations by JWST and Roman may be used to test these scenarios. Moreover, as we explain in section Implications, for a particular DM (sub-component) mass and cross section the formation of Jovian planets is affected as a function of the distance from the GC, enabling a differential detection of the heating signal.
In this work, we have estimated the sensitivity of protoplanets to DM injections using a semi-analytical calculation, with input from simulations in the absence of these injections. Future work should focus on the embedding of DM heat in planetary formation simulations, such that more realistic projections can be obtained. Here we have focused on the core-accretion gas-capture theory of giant planet formation. This is the simplest scenario for the formation of Jovian planets. Alternative scenarios include dynamical processes and migration within the stellar system. Since we focus on the final stages of formation, we expect such alternative scenarios to be qualitatively similar.
@@ Appendix A Dark matter evaporation
The condition (8) relies on the fact that hydrogen evaporates for temperatures above $T_{\rm max}$. Therefore, one would naively say that DM candidates with masses below the mass of hydrogen would evaporate at these temperatures too. However, unlike hydrogen, as we demonstrate below the DM particles are concentrated in the core of the objects. Thus, in the short mean free path regime the lower number density of DM particles in the envelope implies a lower flux of escaping particles (Disrupting Jovian Formation). Here we estimate the evaporation rate of DM particles of a given mass, and compare it to the capture rate.
The DM number density within the protoplanet is obtained by integrating the radial equation Eq. (11) form Ref. Leane and Smirnov (2022), under the assumption that the envelope has constant density $\rho_{E}$ (at $t=\rm Myr$, $\sim 3\times 10^{-8}\rm g\,cm^{-3}$) and the core has a constant density $\rho_{\rm core}$ (at $t=\rm Myr$, $\sim 2\rm g\,cm^{-3}$). We distinguish two regimes, the isothermal regime, where the DM mean free path $\lambda$ in the object is of the order of the object, and DM thermalises with the core at a constant temperature. The second regime where local thermal equilibrium (LTE) is applicable when the DM mean free path is smaller than the size of the object. Thus we have the total DM profile:
@@
$+\theta\left(\frac{R_{\rm core}}{R_{E}}-r\right)\,,$ (19)
where $\theta$ indicates the Heaviside function. For the regime where the mean free path of DM becomes comparable to SM particles the above equation provides us with the DM density within the last scattering surface, to which we apply the Jeans’ escape criterion, demanding that the total surface evaporation rate is below the capture rate. While for the isothermal, and transition regime we use the formalism of Ref. Gould (1987, 1990) to determine the evaporation mass comparing the evaporation rate and the time-scale needed to reach annihilation equilibrium assuming a core temperature of $3000$ K. This leads to the evaporation contour, shown by the dotted lines in Figs. 23. Note that this line is only marking to lowest mass of the DM retained in the object if DM interacts through contact interactions only. In the presence of attractive long range forces, the evaporation contours are significantly moved to lower DM masses, as discussed in Ref. Acevedo et al. (2023).

### v3

@@ Abstract
In Figs.2 and 3 we show constraints on the spin-dependent and spin-independent scattering cross sections respectively from the fact that the formation of Jupiter in our solar system has not stopped at a mass of $30\rm M_{\oplus}$ as purple contours. For comparison we show current direct detection constraints as gray contours. Here we show the maximal reach in the sensitivity to the direct nuclear recoil process, which are least model dependent. The reach is dominated by the CRESST results at lower masses for spin-independent CRESST:2022lqw, and spin-dependent CRESST:2022dtl cross sections, while at higher masses the LUX LUX:2017ree results dominate for spin-dependent interactions, while the XENON experiment dominates the spin-independent sensitivity XENONCollaboration:2023orw. For large spin-independent cross sections, we also show the XQC constraint Erickcek:2007jv; Mahdawi:2018euy. Note that searches based on the Migdal effect CDEX:2022kcd could potentially reach lower DM masses, however, the results have large theoretical uncertainties Cox:2022ekg, and are not displayed here.
@@
The sensitivity ceiling of direct detection experiments at large cross sections for light DM from Ref. Cappiello:2023hza and for heavy DM from Refs. Kavanagh:2017cru; Cappiello:2020lbk is shown. Since at the large cross sections considered the scaling with atomic mass number is uncertain Digman:2019wdm, we do not resale the results for spin-dependent interactions and a more detailed analysis to obtain a precise ceiling values is needed. Note that in contrast the sensitivity of our constraints and suggested search is only affected at cross sections above $10^{-22}\text{cm}^{2}$, where large drift times would require a more detailed analysis of the DM distribution Acevedo:2020gro; Leane:2022hkk and annihilation outside the core.
@@
We do not show complementary cosmology-dependent constraints such as those from the CMB and Lymann-$\alpha$. DM which scatters with a velocity-independent cross-section can also be constrained from the lack of MW satellites produced in such models Nadler:2019zrb. However, as the DM scattering rates could be different in the early universe Elor:2021swj, those bounds are not as universal as the shown late-time sensitivity. Finally, in the case of DM that comprises only a sub-fraction of the total DM abundance, the direct detection constraints are essentially insensitive below the GeV mass scale McKeen:2022poo (and the sensitivity of other constraints also shrinks significantly Li:2022idr), while our search covers significant parameter space even at a DM fraction as low as $f_{\rm DM}\sim 10^{-3}$, shown as solid chartreuse lines.
In Figs.2 and 3, we show the sensitivity to DM interactions from observations of Jovians near the GC, for which we have assumed $\rho_{\rm DM}=100\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=10\rm\,km\,s^{-1}$, where at small cross sections we find that capture by the core is more efficient than capture by the envelope in both the spin-dependent and the spin-independent case. During the period of formation we are interested in, the core-grow is not appreciable, such that $f_{\rm cap}$ remains approximately constant. As the temperature increase needed to satisfy (2) does grow as a function of time, we expect that formation of Jovian planets in the GC halts at $10\rm M_{\oplus}$ in the parameter range indicated by the dashed contours, and proceeds as normal below these contours. Note that the premature ending is already in tension with inferences of Jovian planets in the GC from lensing events Sahu:2006ex; MOA:2009dkq; koshimoto2021no.
Figure 4: Benchmark for spin-dependent dark matter: $m_{\chi}=5$ GeV, $\sigma_{p}=2\times 10^{-28}\,\rm cm^{2}$. The continuous lines show the capture rates at the time when the planet is $10\,\rm M_{\oplus}$ and the dashed lines show the rates at $30\,\rm M_{\oplus}$.
In Fig. 4 we show the capture efficiency as a function of distance from the GC assuming an NFW profile and a Maxwellian velocity dispersion. For small DM scattering cross sections, core capture dominates, preventing any gas accumulation for Jovian planets at some distance from the GC. For larger DM cross sections, envelope capture dominates, leading to the formation of Jovian planets with a decreasing mass the closer to the GC they are formed. This distinct phenomenology can be targeted by future observations of exoplanets.
@@ Discussion
In this letter, we have studied the inhibition of the formation of Jupiter in our solar system and Jovian planets in the Milky Way galaxy by the capture and annihilation of DM. We have shown that the heat generated by DM annihilation can increase the temperature of the protoplanet above the temperature at which the gas mixture evaporates, and determined the corresponding constraints on the parameter space of spin-(in)dependent DM. These constraints are competitive for relatively strongly interacting DM candidates, above the “ceiling” of direct detection constraints. We have shown that these constraints are not significantly weakened if only a subfraction of DM is responsible for the heating of protoplanets.
We have also estimated the projected sensitivities from the existence of Jovian planets in the GC, where the DM density is higher and the typical velocity dispersion lower. Future expolanet observations by JWST and Roman may be used to test these scenarios. Moreover, as we explained in the previous section, for a particular DM (sub-component) mass and cross section the formation of Jovian planets is affected as a function of the distance from the GC, enabling a differential detection of the heating signal.
Here we have focused on the core-accretion gas-capture theory of giant planet formation. This is the simplest scenario for the formation of Jovian planets. Alternative scenarios include dynamical processes and migration within the stellar system. Since we focus on the final stages of formation, we expect such alternative scenarios to be qualitatively similar.
@@ Appendix A Dark matter evaporation
The condition (3) relies on the fact that hydrogen evaporates for temperatures above $T_{\rm max}$. Therefore, one would naively say that DM candidates with masses below the mass of hydrogen would evaporate at these temperatures too. However, unlike hydrogen, as we demonstrate below the DM particles are concentrated in the core of the objects. Thus, in the short mean free path regime the lower number density of DM particles in the envelope implies a lower flux of escaping particles (2). Here we estimate the evaporation rate of DM particles of a given mass, and compare it to the capture rate.
The DM number density within the protoplanet is obtained by integrating the radial equation Eq. (11) form Ref. Leane:2022hkk, under the assumption that the envelope has constant density $\rho_{E}$ (at $t=\rm Myr$, $\sim 3\times 10^{-8}\rm g\,cm^{-3}$) and the core has a constant density $\rho_{\rm core}$ (at $t=\rm Myr$, $\sim 2\rm g\,cm^{-3}$). We distinguish two regimes, the isothermal regime, where the DM mean free path $\lambda$ in the object is of the order of the object, and DM thermalises with the core at a constant temperature. The second regime where local thermal equilibrium (LTE) is applicable when the DM mean free path is smaller than the size of the object. Thus we have the total DM profile:
@@
$+\theta\left(\frac{R_{\rm core}}{R_{E}}-r\right)\,.$ (13)
For the LTE regime the above equation provides us with the DM density within the last scattering surface, to which we apply the Jeans’ escape criterion, demanding that the total surface evaporation rate is below the capture rate. While for the isothermal, and transition regime we use the formalism of Ref. 1987ApJ...321..560G; 1990ApJ...356..302G, which has been applied to various celestail bodies in Ref. Garani:2021feo, to determine the evaporation mass assuming a core temperature of $3000$ K. This leads to the evaporation contour, shown by the dotted lines in Figs. 23. Note that this line is only marking to lowest mass of the DM retained in the object if DM interacts through contact interactions only. In the presence of long range attractive forces, the evaporation contours are significantly moved to lower DM masses, as discussed in Ref. Acevedo:2023owd.